I'm having difficulty solving this quadratic equation. Please help me!
1/5x - 1/10x = 1
My answer came to x=1. I know its wrong. Please help me

Found 2 solutions by rothauserc, MathTherapy:
Answer by rothauserc(4718)

1/5x - 1/10x = 1
:
Note assume x not = 0
:
multiply both sides of = by 10x
:
2 - 1 = 10x
:
10x = 1
:
**************
x = 1/10 = 0.1
**************
:
check answer
:
(1 / (5 * 0.1)) - (1 / (10 * 0.1)) = 1
:
10/5 - (1/1) = 1
:
2 - 1 = 1
:
1 = 1
:
answer checks
:

Answer by MathTherapy(10551)


I'm having difficulty solving this quadratic equation. Please help me!
1/5x - 1/10x = 1
My answer came to x=1. I know its wrong. Please help me
First off, you have to use PARENTHESES, so one can KNOW whether the x is part of the denominator, or numerator.
In other words, state whether it's: